write the equation -4x + 2y = 1 in slope-intercept form. give the slope and the y-intercept
Answer by mananth(16946)

-4x + 2y = 1
you have to write in the form
y = mx+b
2y=4x+1
/2
y = 2x +1/2
slope = 2
and y intercept = 1/2